Jeans are a new fashion among adults. Below given data provide the fraction of Indian
men of age 20 years or older having five or more wearable jeans along with the fraction
of the Indian population of these age groups. Find the unconditional probability of an
adult 20 years or older having five or more wearable jeans.
Groups & Ages
20-24 25-34 35-49 50-64 above 65
Fraction with 5 or more jeans .26 .20 .13 .18 .14
Fraction adults 20 or older .09 .20 .31 .23 .17
1
Expert's answer
2020-08-03T18:35:12-0400
"\begin{matrix}
 & Fraction\ with\ 5 & Fraction \ adults \\
 & or\ more\ jeans & 20\ or\ older \\
 20-24 & 0.26 & 0.09 \\
25-34 & 0.20 & 0.20 \\
35-49 & 0.13 & 0.31 \\
50-64 & 0.18 & 0.23 \\
above\ 65 & 0.14 & 0.17
\end{matrix}"

Check


"0.09+0.20+0.31+0.23+0.17=1"

"P(X\geq5)=0.26(0.09)+0.20(0.20)+"

"+0.13(0.31)+0.18(0.23)+0.14(0.17)=0.1689"